$7 trilliorn $1.25 trillion 350 million 50 million What is the GDP per capita for Country B? $ 6250
Consider the table of GDP and population for several imaginary countries. 21.0 191 175.00 Wrigleyville 551.0 2315 Dinky 1.900 Using this information, please answer the questions. If there is not enough information to answer a question, please enter -11. A. What is the GDP per capita of Longhornland in U.S. dollars? Express your answer rounded to one decimal place. S 16.2 B. What is the GDP of Dinkytown in millions of U.S. dollars? Express your answer rounded to one decimal place. S 221.1 0
Answer #1

1.Per Capita GDP for country A =Gross Domestic Product/Population

So, Per Capita GDP for country A = $7,000,000,000,000/350,000,000 = $20,000

2.Per Capita GDP for country A =Gross Domestic Product/Population

So, Per Capita GDP for country A = $1,25,000,000,000,000/50,000,000 = $25,000

3. Country B has largest average GDP per capita.

4. GDP Per capita of Longhornland = Gross Domestic Product/Population

= $23,15,000,000/19,10,00,000 = 12.12 US$.

It shows that the economic production and growth of a nation and is one of the primary indicators used to determine the overall well-being of a country’s economy and standard of living. The economic policymakers should assess whether the economy is weakening or progressing if it needs improvements or restrictions, and if threats of recession or inflation are imminent.

5. GDP of Dinkytown is 332.5 US$.